(1 point) An oceanographic profiling instrument was pressure tested up to 15, 000 psig. After reviewing the test chamber pressure data, you notice a slight drop in pressure at a test pressure of 7000 psig, which you surmise is due to the test chamber fluid leaking into the instrument. What is the maximum oceanwater depth above which this instrument can safely operate? At significant depths, the density of ocean water is approximately 1028 kg/mi3. km
Answer #1

Instrument can work safely up to 7000 psig
We know
1 psig = 0.0689 bar
700 psig = 482.633 bar
so safe pressure (P) = 482.633 *105 N/m2
Now we know that the pressure at depth d will be
P = ho gd
where ho is density of ocean water = 1028 kg/m3
g is acceleration due to gravity = 9.81 m/s2
d is depth of water
P (1028) 9.81d
We know that the safe pressure (P) = 482.633 *105 N/m2
482.633 * 105-(1028) * 9.81d
d = 4.786*103 m
d = 4.786 Km
Hence instrument can work safely up to 4.786 Km depth